OpenStudy (anonymous):

5x+4y-5z=26 2x-5y+2z=13 4x-3y+4z=19

OpenStudy (anonymous):

2x-5y+2z=13 4x-3y+4z=19 u can eeliminate x and z

OpenStudy (anonymous):

are you solving this with a matrix?

OpenStudy (anonymous):

no substitoution

OpenStudy (anonymous):

come on amanda you can get y from these 2 eqs 2x-5y+2z=13 4x-3y+4z=19

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k next step ten minutes.

OpenStudy (anonymous):

10 mins? it will take u at least 3 mins to solve it here is x x = 5 solve for z

OpenStudy (anonymous):

z=5

OpenStudy (anonymous):

i mean 4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

5x+4y-5z=26 5*5-4 -5z =26 21 - 5z = 26 -5z = 5 z = -1

OpenStudy (anonymous):

next question, u need to get the A

OpenStudy (anonymous):

this one is really long.